Contirm your decision usi t 6Which Design should be chosen from the designs listed below? A 10 year study period and MARR-10% per year is to be used. No market value considered Capital investment Annual receipts Annual expenses Design A $170,000 114,000 70,000 Design BDesign C S300,000 130,000 $330,000 147,000 79,000 64,000
Answer #1

To select the best investment we will use annual worth method.

AW = - Initial investment (A/P,I,n) + receipts - expenses

n = 10 years

I = 10% = 0.1

(A/P,I,n) = I (1+I)n / ((1+I)n​​​​​​ - 1)

(A/P,10%,10) = 0.1(1+0.1)10/( (1+0.1)10-1) = 0.163

Design investment (A/P,10%,10) (-)investment X (A/P,10%,10) (-)Annual expenses Annual receipt AW
A 170000 0.163 27710 70000 114000 16290
B 330000 0.163 53790 79000 147000 14120
C 300000 0.163 48900 64000 130000 17100

According to Annual worth method I will choose deign C